# Quantized tensor networks for solving the Vlasov-Maxwell equations

### v3

@@ Abstract
The Vlasov-Maxwell equations provide an ab-initio description of collisionless plasmas, but solving them is often impractical because of the wide range of spatial and temporal scales that must be resolved and the high dimensionality of the problem. In this work, we present a quantum-inspired semi-implicit Vlasov-Maxwell solver that utilizes the quantized tensor network (QTN) framework. With this QTN solver, the cost of grid-based numerical simulation of size $N$ is reduced from $\mathcal{O}(N)$ to $\mathcal{O}(\text{poly}(D))$, where $D$ is the “rank” or “bond dimension” of the QTN and is typically set to be much smaller than $N$. We find that for the five-dimensional test problems considered here, a modest $D=64$ appears to be sufficient for capturing the expected physics despite the simulations using a total of $N=2^{36}$ grid points, which would require $D=2^{18}$ for full-rank calculations. Additionally, we observe that a QTN time evolution scheme based on the Dirac-Frenkel variational principle allows one to use larger time steps than prescribed by the Courant-Friedrichs-Lewy (CFL) constraint. As such, this work demonstrates that the QTN format is a promising means of approximately solving the Vlasov-Maxwell equations with significantly reduced cost.
@@ 1 Introduction
This concept was first introduced several years ago, and is referred to as quantics or quantized tensor trains (QTTs) because the data is artificially folded into a high-dimensional tensor representing many small “quants” of data (Khoromskij, 2011; Dolgov et al., 2012; Oseledets, 2009, 2010). (Tensor trains are tensor networks with a 1-D geometry.) It was also proven that several one-dimensional functions (exponential, trigonometric, polynomial) and common operators (e.g., differential operators) on a uniform grid can be efficiently represented in the QTT format (Khoromskij, 2011; Kazeev & Khoromskij, 2012; Kazeev et al., 2013a). It follows that one can often obtain a loose bound on the QTT rank for other smooth functions by expanding them in terms of exponential or polynomial basis sets (Dolgov et al., 2012), though tighter bounds can be obtained (Lindsey, 2023). However, non-smooth functions can have efficient representations too; the Heaviside step function can be represented with just rank two, triangle functions with rank three (Gourianov et al., 2022). The underlying intuition is that functions with some degree of scale locality, meaning that features on long length scales are relatively independent of behavior at small length scales, can be represented as a tensor network with low rank. Since it is believed that turbulent systems also exhibit some degree of scale locality, tensor networks may even be able to efficiently capture turbulent dynamics (Gourianov, 2022).
@@ 2 Quantized Tensor Networks
Quantized tensor networks are a multi-scale ansatz built from a set of interconnected high-dimensional matrices typically referred to as tensors (Oseledets, 2009; Khoromskij, 2011; Dolgov, 2014). For more background information on QTNs, we refer readers to the previous references as well as Ripoll (2021) and Ye & Loureiro (2022).
Suppose a 1-D function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h}(x)$ is discretized into $N=d^{L}$ grid points. One can quantize the representation, by which one defines artificial dimensions to describe the function. This effectively appears as reshaping the vector into an $L$-dimensional tensor with size $d$ along each dimension, $h(x)\cong h(i_{1},i_{2},\ldots,i_{L})$, where the indices $i_{1},i_{2},...i_{L}\in\{0,\ldots,d-1\}$ encode the grid points in $x$. For example, if $d=2$, one possible encoding is a binary mapping, such that the $i^{\text{th}}$ grid point is indexed by its binary string $(i_{1},i_{2},...,i_{L})_{2}$. In this case, the indices sequentially correspond to finer grid resolutions. Another possible encoding is a signed register, in which one takes the two’s complement of the physical indices conditioned on the first index, such that $i\cong(i_{1},i_{1}+i_{2},...,i_{1}+i_{L})_{2}$. This mirror-like mapping may be more efficient for symmetric functions, since the second half of the data is flipped with respect to the original (Ripoll, 2021). It is used later in Section 5 when representing Gaussian functions in velocity space.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h}(i_{1},i_{2},...,i_{L})=\sum_{\alpha_{1}=1}^{r_{1}}...\sum_{\alpha_{L-1}=1}^{r_{L-1}}M^{(1)}_{\alpha_{1}}(i_{1})M^{(2)}_{\alpha_{1},\alpha_{2}}(i_{2})\ldots M^{(L)}_{\alpha_{L-1}}(i_{L}),$ (1)
where $M^{(p)}$ is the $p^{\text{th}}$ tensor in the chain with one “physical” bond ($i_{p}$) of size $d$ that is used to access the actual data, and two “virtual” bonds ($\alpha_{p-1},\alpha_{p}$) of size $r_{p-1}$ and $r_{p}$ that connect the tensor to its two neighbors. Upon contraction of the tensor network, the sums in the above equation are performed and the virtual bonds no longer exist. (The exceptions are $M^{(1)}$, which only has one virtual bond ($\alpha_{1}$) of size $r_{1}$, and $M^{(L)}$, which also only has one virtual bond ($\alpha_{L-1}$) of size $r_{L-1}$.) The simplest way to obtain the tensors in Eq. (1) is to perform a series of singular value decompositions (Oseledets, 2010), isolating each dimension $i_{p}$ one at a time. If no approximation is made and the vector is full-rank, then rank $r_{p}$ is $\text{min}(d^{p},d^{L-p})$. The rank is largest at the middle of the tensor train, with $r_{L/2}=d^{L/2}=\sqrt{N}$. To obtain a good low-rank approximation, one would keep only the $D$ largest singular values at each decomposition step. This parameter is referred to as the bond dimension or rank of the tensor train. (The terms bond dimension and rank will be used interchangeably.) If the original dataset can be represented as a QTT of small bond dimension (ideally $D\sim\log(N)$) with only some small tolerable amount of error, the dataset is described as compressible and can be efficiently represented as a QTT. For concrete examples of common functions that can be efficiently represented in the QTT format, we refer the reader to Dolgov (2014, p. 56).
@@
$O(x^{\prime},x)$ $\cong O(o_{1},o_{2},...,o_{L};,i_{1},i_{2},...,i_{L})$
$=\sum_{\alpha_{1}=1}^{r_{1}}...\sum_{\alpha_{L-1}=1}^{r_{L-1}}M^{(1)}_{\alpha_{1}}(o_{1},i_{1})M^{(2)}_{\alpha_{1},\alpha_{2}}(o_{2},i_{2})\ldots\ldots M^{(L)}_{\alpha_{L-1}}(o_{L},i_{L})$ (2)
where each tensor $M^{(p)}$ now has two physical bonds ($o_{p},i_{p}$) that are used to access the output $x^{\prime}$ and input $x$ of the discretized operator $O$.
The QTT representation can be considered “quantum-inspired” since mapping classical data onto a quantum computer could be done in a similar fashion (Ripoll, 2021; Lubasch et al., 2018, 2020; Jaksch et al., 2023). In this case, each tensor would correspond to one qubit. The physical bond in the tensor represents the state of the qubit, while the virtual bonds capture the entanglement between the qubits. The QTT-Operators are analogs to quantum gates or quantum circuits applied to the quantum computer.
@@
$h(x_{1,i},$ $x_{2,i},\dots,x_{K,i})$
$\cong h(i^{(1)}_{1},\dots,i^{(1)}_{L},i^{(2)}_{1},\dots,i^{(2)}_{L},\dots,i^{(K)}_{1},\dots,i^{(K)}_{L})$
$=\sum_{\gamma_{1}=1}^{r_{1}}...\sum_{\gamma_{K-1}=1}^{r_{K-1}}B^{(1)}_{\gamma_{1}}(i_{1},...i_{L})B^{(2)}_{\gamma_{1},\gamma_{2}}(j_{1},...,j_{L})\ldots\ldots B^{(K)}_{\gamma_{L-1}}(i^{(K)}_{1},...,i^{(K)})\,,$ (3)
where $B^{(k)}$ is a tensor train for the $k^{\text{th}}$ dimension
$B^{(k)}_{\gamma_{k},\gamma_{k+1}}(i_{1},...,i_{L})=$
$\hskip 14.22636pt\sum_{\beta_{k}}\tilde{M}^{(k)}_{\gamma_{k},\gamma_{k+1},\beta_{k}}\biggl{(}\sum_{\alpha_{k,1}}...\sum_{\alpha_{k,L-1}}M^{(k,1)}_{\alpha_{k},\alpha_{(k,1)}}(i^{(k)}_{1})M^{(k,2)}_{\alpha_{(k,1)},\alpha_{(k,2)}}(i^{(k)}_{2})\ldots M^{(k,L)}_{\alpha_{(k,L-1)}}(i^{(k)}_{L})\biggr{)}\,.$ (4)
To simplify notation, we assumed that all problems dimensions have the same number of grid points. However, this is not a necessary requirement. In Eq. (4), tensor $\tilde{M}^{(k)}$ is the $k^{\text{th}}$ tensor in the spine of the comb. Note that it does not have a physical index. In the low-rank approximation, the size of each dimension is capped at $S$. For simplicity, this work uses the same maximum bond dimension for all virtual bonds (i.e., $S=D$), though it may be advantageous to use different bond dimensions for the bonds in the branches ($\alpha_{(k,i)}$), along the spine ($\gamma_{k}$), and between the spine to the branches ($\beta_{k}$).
@@ 4 Vlasov-Maxwell Solver
We implement a semi-implicit finite-difference solver of the Vlasov-Maxwell equations utilizing the quantized tensor network (QTN) format. The distribution functions for each species and each component of the electromagnetic fields are represented as QTNs with bond sizes of at most $D$. Likewise, the operators acting on these fields (e.g., the derivatives) are also represented in the QTN format (see Ye & Loureiro (2022) for details).
@@ 4.2 Semi-Lagrangian time evolution
the operator $A_{SL}$ is effectively 2-dimensional, as it can be written in the form $A_{SL}=\tilde{A}_{x,v_{x}}\otimes\mathbb{I}_{y,z,v_{y},v_{z}}$ (the subscripts denote the problem dimensions on which the operator acts). It is observed that the rank of the QTN representation of $\tilde{A}_{x,v_{x}}$ depends on the largest magnitude of integer $p$; so long as moderate time step and grid resolution is used the bond dimension remains manageable. Furthermore, if the time step is held constant, the QTN decomposition only needs to be performed once.
Once in the QTT format, the matrix-vector multiplication is performed using the density-matrix method (McCulloch, 2007). Details of this algorithm for the comb geometry are given in Section of the Supplementary Information (SI), and the computational costs are listed in Table 1. Since the bond dimension of the operator $A_{SL}$ is expected to be modest for all time, its contribution to the scaling can be neglected, yielding costs that like $\mathcal{O}(D^{3})$ for the QTT geometry and $\mathcal{O}(D^{4})$ for the QTC geometry. For the Orszag-Tang vortex calculation, the observed scaling is closer to $\mathcal{O}(D^{2.5})$ (see Section VI.A.4 of the SI).
@@ 4.3 Time-dependent variational principle
We can instead consider a local time evolution scheme, in which individual tensors in the QTN are propagated forward in time in a sequential fashion. In quantum simulation, one such algorithm is the time-dependent variational principle (TDVP) scheme (Haegeman et al., 2016). This scheme is also closely related to the dynamical low-rank algorithm (Khoromskij et al., 2012; Einkemmer & Lubich, 2018), which has been used for solving PDEs in the non-quantized tensor train community. More recently, the same underlying principle has been used in a quantum algorithm for solving the Schrodinger-Poisson equation (Cappelli et al., 2023).
@@
TDVP utilizes the Dirac-Frenkel variational principle, which states that the dynamics of a unitary system at any given point in time lies in the tangent space of the current state (Frenkel, 1934; Lasser & Su, 2022). By treating each tensor in the QTN as a variational parameter, one can obtain the effective dynamics for each tensor (given by anti-Hermitian operator $A_{\text{eff}}^{(i)}$ for the $i^{\text{th}}$ tensor) and sequentially evolve them forward in time. More specifically, if the unitary dynamics acting on state $\psi$ is described by anti-Hermitian operator $A$, the equation of motion for the $i^{\text{th}}$ tensor in the QTN $M_{i}$ is
$\frac{\partial}{\partial t}M_{i}$ $=A_{\text{eff}}^{(i)}M_{i}\,,$ (11)
@@
$A_{\text{eff}}^{(i)}=\partial_{M_{i}}\,\partial_{M_{i}^{*}}\langle\psi(M_{1},...,M_{L})\,|\,A\,|\,\psi(M_{1},...M_{L})\rangle\,,$ (12)
assuming that the QTN is canonicalized such that the orthogonality center is located at the $i^{\text{th}}$ tensor. As a result, the original dynamics is decomposed into roughly $2KL$ equations of motion of size $\mathcal{O}(D^{2})$. (The factor of two arises from the algorithm when sweeping across all tensors.) We refer the reader to the Section IV of the SI and Paeckel et al. (2019) for a more in-depth explanation and the algorithmic details.
In this work, Eq. (11) is solved using RK4, so the computational cost scales like $\mathcal{O}(D^{3}D_{W}d+D^{2}D_{W}^{2}d^{2})$, where $D_{W}$ is the bond dimension of the QTT-O representation for time evolution operator $A$. For the QTC geometry, treatment of the tensors along the spine formally scales like $\mathcal{O}(S^{2}S_{W}^{3}+S^{4}S_{W}^{2})$, where $S$ and $S_{W}$ are the bond dimensions of the spine tensors in the QTC representations of state $\psi$ and operator $A$. In the case of the Vlasov equation with two spatial dimensions, $S_{W}$ can be neglected, but $D_{W}\sim D$ due to the nonlinear term. Setting $S=D$, the overall expected scaling is $\mathcal{O}(D^{4})$, which is indeed roughly the scaling observed for the Orszag-Tang vortex calculation (see Section VI.A.4 of the SI).
@@
Particularly interesting is that even when Eq. (11) is solved approximately using RK4, we observe that the local methods allow one to use larger time steps than what global methods are limited to (e.g., by the Courant-Friedrich-Levy (CFL) constraint (Courant et al., 1967)). In the following example, it appears that the time steps of the local algorithm can be larger by at least a factor of four. This is perhaps possible because global information is incorporated into the effective operator for each individual tensor. The ability to bypass CFL is crucial for QTT finite difference solvers, as it allows one to use higher spatial resolution (for improved accuracy) without requiring one to significantly increase temporal resolution.

### v4

@@ Abstract
The Vlasov-Maxwell equations provide an ab-initio description of collisionless plasmas, but solving them is often impractical because of the wide range of spatial and temporal scales that must be resolved and the high dimensionality of the problem. In this work, we present a quantum-inspired semi-implicit Vlasov-Maxwell solver that utilizes the quantized tensor network (QTN) framework. With this QTN solver, the cost of grid-based numerical simulation of size $N$ is reduced from $\mathcal{O}(N)$ to $\mathcal{O}(\text{poly}(D))$, where $D$ is the “rank” or “bond dimension” of the QTN and is typically set to be much smaller than $N$. We find that for the five-dimensional test problems considered here, a modest $D=64$ appears to be sufficient for capturing the expected physics despite the simulations using a total of $N=2^{36}$ grid points, which would require $D=2^{18}$ for full-rank calculations. Additionally, we observe that a QTN time evolution scheme based on the Dirac-Frenkel variational principle allows one to use somewhat larger time steps than prescribed by the Courant-Friedrichs-Lewy (CFL) constraint. As such, this work demonstrates that the QTN format is a promising means of approximately solving the Vlasov-Maxwell equations with significantly reduced cost.
@@ 1 Introduction
This concept was first introduced several years ago, and is referred to as quantics or quantized tensor trains (QTTs) because the data is artificially folded into a high-dimensional tensor representing many small “quants” of data (Khoromskij, 2011; Dolgov et al., 2012; Oseledets, 2009, 2010). (Tensor trains are tensor networks with a 1-D geometry.) It was also proven that several one-dimensional functions (exponential, trigonometric, polynomial) and common operators (e.g., differential operators) on a uniform grid can be efficiently represented in the QTT format (Khoromskij, 2011; Kazeev & Khoromskij, 2012; Kazeev et al., 2013b). It follows that one can often obtain a loose bound on the QTT rank for other smooth functions by expanding them in terms of exponential or polynomial basis sets (Dolgov et al., 2012), though tighter bounds can be obtained (Lindsey, 2023). However, non-smooth functions can have efficient representations too; the Heaviside step function can be represented with just rank two, triangle functions with rank three (Gourianov et al., 2022). The underlying intuition is that functions with some degree of scale locality, meaning that features on long length scales are relatively independent of behavior at small length scales, can be represented as a tensor network with low rank. Since it is believed that turbulent systems also exhibit some degree of scale locality, tensor networks may even be able to efficiently capture turbulent dynamics (Gourianov, 2022).
@@ 2 Quantized Tensor Networks
Quantized tensor networks are a multi-scale ansatz built from a set of interconnected high-dimensional matrices typically referred to as tensors (Oseledets, 2009; Khoromskij, 2011; Dolgov, 2014). For more background information on QTNs, we refer readers to the previous references as well as Garcia-Ripoll (2021) and Ye & Loureiro (2022).
Suppose a 1-D function ${h}(x)$ is discretized into $N=d^{L}$ grid points. One can quantize the representation, by which one defines artificial dimensions to describe the function. This effectively appears as reshaping the vector into an $L$-dimensional tensor with size $d$ along each dimension, $h(x)\cong\textsf{h}(i_{1},i_{2},\ldots,i_{L})$, where the indices $i_{1},i_{2},...i_{L}\in\{0,\ldots,d-1\}$ encode the grid points in $x$. For example, if $d=2$, one possible encoding is a binary mapping, such that the $i^{\text{th}}$ grid point is indexed by its binary string $(i_{1},i_{2},...,i_{L})_{2}$. In this case, the indices sequentially correspond to finer grid resolutions. Another possible encoding is a signed register, in which one takes the two’s complement of the physical indices conditioned on the first index, such that $i\cong(i_{1},i_{1}+i_{2},...,i_{1}+i_{L})_{2}$. This mirror-like mapping may be more efficient for symmetric functions, since the second half of the data is flipped with respect to the original (Garcia-Ripoll, 2021). It is used later in Section 5 when representing Gaussian functions.
@@
$\textsf{h}(i_{1},i_{2},...,i_{L})=\sum_{\alpha_{1}=1}^{r_{1}}...\sum_{\alpha_{L-1}=1}^{r_{L-1}}\textsf{M}^{(1)}_{\alpha_{1}}(i_{1})\textsf{M}^{(2)}_{\alpha_{1},\alpha_{2}}(i_{2})\ldots\textsf{M}^{(L)}_{\alpha_{L-1}}(i_{L}),$ (1)
where $\textsf{M}^{(p)}$ is the $p^{\text{th}}$ tensor in the chain with one “physical” bond ($i_{p}$) of size $d$ that is used to access the actual data, and two “virtual” bonds ($\alpha_{p-1},\alpha_{p}$) of size $r_{p-1}$ and $r_{p}$ that connect the tensor to its two neighbors. Upon contraction of the tensor network, the sums in the above equation are performed and the virtual bonds no longer exist. (The exceptions are $\textsf{M}^{(1)}$, which only has one virtual bond ($\alpha_{1}$) of size $r_{1}$, and $\textsf{M}^{(L)}$, which also only has one virtual bond ($\alpha_{L-1}$) of size $r_{L-1}$.) The simplest way to obtain the tensors in Eq. (1) is to perform a series of singular value decompositions (Oseledets, 2010), isolating each dimension $i_{p}$ one at a time. If no approximation is made and the vector is full-rank, then rank $r_{p}$ is $\text{min}(d^{p},d^{L-p})$. The rank is largest at the middle of the tensor train, with $r_{L/2}=d^{L/2}=\sqrt{N}$. To obtain a good low-rank approximation, one would keep only the $D$ largest singular values at each decomposition step. This parameter is referred to as the bond dimension or rank of the tensor train. (The terms bond dimension and rank will be used interchangeably.) If the original dataset can be represented as a QTT of small bond dimension (ideally $D\sim\log(N)$) with only some small tolerable amount of error, the dataset is described as compressible and can be efficiently represented as a QTT. For concrete examples of common functions that can be efficiently represented in the QTT format, we refer the reader to Dolgov (2014, p. 56).
@@
$\textsf{O}(x^{\prime},x)$ $\cong\textsf{O}(o_{1},o_{2},...,o_{L};i_{1},i_{2},...,i_{L})$
$=\sum_{\alpha_{1}=1}^{r_{1}}...\sum_{\alpha_{L-1}=1}^{r_{L-1}}\textsf{M}^{(1)}_{\alpha_{1}}(o_{1},i_{1})\textsf{M}^{(2)}_{\alpha_{1},\alpha_{2}}(o_{2},i_{2})\ldots\textsf{M}^{(L)}_{\alpha_{L-1}}(o_{L},i_{L})$ (2)
where each tensor $\textsf{M}^{(p)}$ now has two physical bonds ($o_{p},i_{p}$) that are used to access the output $x^{\prime}$ and input $x$ of the discretized operator O.
The QTT representation can be considered “quantum-inspired” since mapping classical data onto a quantum computer could be done in a similar fashion (Garcia-Ripoll, 2021; Lubasch et al., 2018, 2020; Jaksch et al., 2023). In this case, each tensor would correspond to one qubit. The physical bond in the tensor represents the state of the qubit, while the virtual bonds capture the entanglement between the qubits. The QTT-Operators are analogs to quantum gates or quantum circuits applied to the quantum computer.
@@
$\textsf{h}(x_{1,i},$ $x_{2,i},\dots,x_{K,i})$
$\cong\textsf{h}(i^{(1)}_{1},\dots,i^{(1)}_{L},i^{(2)}_{1},\dots,i^{(2)}_{L},\dots,i^{(K)}_{1},\dots,i^{(K)}_{L})$
$=\sum_{\gamma_{1}=1}^{r_{1}}...\sum_{\gamma_{K-1}=1}^{r_{K-1}}\textsf{B}^{(1)}_{\gamma_{1}}(i^{(1)}_{1},...,i^{(1)}_{L})\textsf{B}^{(2)}_{\gamma_{1},\gamma_{2}}(i^{(2)}_{1},...,i^{(2)}_{L})\ldots\textsf{B}^{(K)}_{\gamma_{L-1}}(i^{(K)}_{1},...,i^{(K)})\,,$ (3)
where $\textsf{B}^{(k)}$ is a tensor train for the $k^{\text{th}}$ problem dimension
$\textsf{B}^{(k)}_{\gamma_{k},\gamma_{k+1}}(i^{(k)}_{1},...,i^{(k)}_{L})=$
$\hskip 14.22636pt\sum_{\beta_{k}}\tilde{\textsf{M}}^{(k)}_{\gamma_{k},\gamma_{k+1},\beta_{k}}\biggl{(}\sum_{\alpha_{(k,1)}}...\sum_{\alpha_{(k,L-1)}}\textsf{M}^{(k,1)}_{\beta_{k},\alpha_{(k,1)}}(i^{(k)}_{1})\textsf{M}^{(k,2)}_{\alpha_{(k,1)},\alpha_{(k,2)}}(i^{(k)}_{2})\ldots\textsf{M}^{(k,L)}_{\alpha_{(k,L-1)}}(i^{(k)}_{L})\biggr{)}\,.$ (4)
To simplify notation, we assumed that all problems dimensions have the same number of grid points. However, this is not a necessary requirement. In Eq. (4), tensor $\tilde{\textsf{M}}^{(k)}$ is the $k^{\text{th}}$ tensor in the spine of the comb. Note that it does not have a physical index. In the low-rank approximation, the size of each dimension is capped at $S$. For simplicity, this work uses the same maximum bond dimension for all virtual bonds (i.e., $S=D$), though it may be advantageous to use different bond dimensions for the bonds in the branches ($\alpha_{(k,i)}$), along the spine ($\gamma_{k}$), and between the spine to the branches ($\beta_{k}$).
@@ 4 Vlasov-Maxwell Solver
We implement a semi-implicit finite-difference solver of the Vlasov-Maxwell equations utilizing the quantized tensor network (QTN) format. The distribution functions for each species and each component of the electromagnetic fields are represented as QTNs with bond sizes of at most $D$. Likewise, the operators acting on these fields (e.g., the derivatives) are also represented in the QTN format (see Section I of the SI and Ye & Loureiro (2022) for details).
@@ 4.2 Semi-Lagrangian time evolution
the operator $\textsf{A}_{SL}$ is effectively 2-dimensional, as it can be written in the form $\textsf{A}_{SL}=\tilde{\textsf{A}}_{x,v_{x}}\otimes\mathbb{I}_{y,z,v_{y},v_{z}}$ (the subscripts denote the problem dimensions on which the operator acts). It is observed that the rank of the QTN representation of $\tilde{\textsf{A}}_{x,v_{x}}$ depends on the largest magnitude of integer $p$; so long as moderate time step and grid resolution is used the bond dimension remains manageable. Furthermore, if the time step is held constant, the QTN decomposition only needs to be performed once.
Once in the QTT format, the matrix-vector multiplication is performed using the density-matrix method (McCulloch, 2007). Details of this algorithm for the comb geometry are given in Section III of the SI, and the computational costs are listed in Table 1. Since the bond dimension of the operator $\textsf{A}_{SL}$ is expected to be modest for all time, its contribution to the scaling can be neglected, yielding costs that scale like $\mathcal{O}(D^{3})$ for the QTT geometry and $\mathcal{O}(D^{4})$ for the QTC geometry. For the Orszag-Tang vortex calculation, the observed scaling is closer to $\mathcal{O}(D^{2.5})$ (see Section VI.A.4 of the SI).
@@ 4.3 Time-dependent variational principle
We can instead consider a local time evolution scheme, in which individual tensors in the QTN are propagated forward in time in a sequential fashion. In quantum simulation, one such algorithm is the time-dependent variational principle (TDVP) scheme (Haegeman et al., 2016). This scheme is also closely related to the dynamical low-rank algorithm (Khoromskij et al., 2012; Einkemmer & Lubich, 2018), which has been used for solving PDEs in the non-quantized tensor train community. More recently, the same underlying principle has been used in a quantum algorithm for solving the Schrodinger-Poisson equation (Cappelli et al., 2024).
@@
TDVP utilizes the Dirac-Frenkel variational principle, which states that the dynamics of a unitary system at any given point in time lies in the tangent space of the current state (Frenkel, 1934; Lasser & Su, 2022). By treating each tensor in the QTN as a variational parameter, one can obtain the effective dynamics for each tensor (given by anti-Hermitian operator $\textsf{A}_{\text{eff}}^{(i)}$ for the $i^{\text{th}}$ tensor) and sequentially evolve them forward in time. More specifically, if the unitary dynamics acting on state $\psi$ is described by anti-Hermitian operator A, the equation of motion for the $i^{\text{th}}$ tensor in the QTN $\textsf{M}_{i}$ is
$\frac{\partial}{\partial t}\textsf{M}_{i}$ $=\textsf{A}_{\text{eff}}^{(i)}\textsf{M}_{i}\,,$ (11)
@@
$\textsf{A}_{\text{eff}}^{(i)}=\partial_{\textsf{M}_{i}}\,\partial_{\textsf{M}_{i}^{*}}\langle\psi(\textsf{M}_{1},...,\textsf{M}_{L})\,|\,\textsf{A}\,|\,\psi(\textsf{M}_{1},...\textsf{M}_{L})\rangle\,,$ (12)
assuming that the QTN is canonicalized such that the orthogonality center is located at the $i^{\text{th}}$ tensor. As a result, the original dynamics is decomposed into roughly $2KL$ equations of motion of size $\mathcal{O}(D^{2})$. (The factor of two arises from the algorithm when sweeping across all tensors.) For a second-order scheme, one updates all the tensors sequentially in a sweeping fashion with increasing $i$ and then decreasing $i$, using half of the original time step at each update. We refer the reader to the Section IV of the SI and Paeckel et al. (2019) for a more in-depth explanation and the algorithmic details.
In this work, Eq. (11) is solved using RK4, so the computational cost scales like $\mathcal{O}(D^{3}D_{W}d+D^{2}D_{W}^{2}d^{2})$, where $D_{W}$ is the bond dimension of the QTT-O representation for time evolution operator $A$. For the QTC geometry, treatment of the tensors along the spine formally scales like $\mathcal{O}(S^{2}S_{W}^{3}+S^{4}S_{W}^{2})$, where $S$ and $S_{W}$ are the bond dimensions of the spine tensors in the QTC representations of state $\psi$ and operator $A$. In the case of the Vlasov equation with two spatial dimensions, $S_{W}$ can be neglected, but $D_{W}\sim D$ due to the nonlinear term. Setting $S=D$, the overall expected scaling is $\mathcal{O}(D^{4})$, which is what is observed for the Orszag-Tang vortex calculation (see Section VI.A.4 of the SI).
@@
Particularly interesting is that even when Eq. (11) is solved approximately using RK4, we observe that the local methods allow one to use larger time steps than what global methods are limited to (e.g., by the Courant-Friedrich-Levy (CFL) constraint (Courant et al., 1967)). In the following example, it appears that the time steps of the local algorithm can be larger by about a factor of two. This is perhaps possible because global information is incorporated into the effective operator for each individual tensor. The ability to go beyond CFL is crucial for QTN finite difference solvers, as it allows one to use higher spatial resolution (for improved accuracy) without requiring one to significantly increase temporal resolution.